Kubin Companys relevant range of production is 21,000 to 25,000 units. When it produces and sells 23,000 units,its average costs per unit are as follows: Average Cost per Unit Direct materials Direct labor Variable manufacturing overhead Fixed manufacturing overhead Fixed selling expense Fixed administrative expense Sales commissions Variable administrative expense 8.10 s 4.66 $2.10 Required: 1. For financial accounting purposes, what is the total amount of product costs incurred to make 23,000 units? 2. For financial accounting purposes, what is the total amount of period costs incurred to sell 23,000 units? 3. For financial accounting p 4. For financial accounting purposes, what is the total amount of penod costs incurred to sell 21,000 units? s, what is the total amount of product costs incurred to make 25,000 units? (For all requirements, do not round intermediate calculations.) 1. Total amount of product costs 2. Total amount of period costs incurred 3. Total amount of product costs S273,700 S249 of
Answer #1

Solution 1:

Total amount of product costs = (Direct materials + Direct labor + Variable manufacturing overhead + Fixed manufacturing overhead) * Nos of units

= ($8.10 + $5.10 + $2.60 + $6.10) * 23000 = $503,700

Solution 2:

Total amount of period costs incurred to sell 23000 units = (Fixed selling expense + Fixed administrative expense + sales commissions + Variable administrative expense)* Nos of units

= ($4.60 + $3.60 + $2.10 + $1.60) * 23000 = $273,700

Solution 3:

Total amount of product costs at 25000 units = (Direct materials + Direct labor + Variable manufacturing overhead)* Nos of units + Fixed manufacturing overhead

= ($8.10 + $5.10 + $2.60)*25000 + 23000*$6.10 = $535,300

Solution 4:

Total amount of period costs incurred to sell 21000 units = Fixed selling expense + Fixed administrative expense + sales commissions + Variable administrative expense

= (23,000*$4.60) + (23000*$3.60) + (21000*$2.10) + (21000*$1.60)

= $105,800 + $82,800 + $44,100 + $33,600

= $266,300